# Physics-informed and Unsupervised Riemannian Domain Adaptation for Machine Learning on Heterogeneous EEG Datasets

### v1

@@ I Introduction
While machine learning is a promising tool for large-scale EEG analysis, its effectiveness can be compromised by dataset shifts. This problem occurs when a model is applied to unseen data (target) with a different distribution than the data on which it was trained (source) [5]. In EEG data, this variability arises from inter-subject differences, inter-session variability, and differences in recording equipment, e.g., amplifiers and electrode nets. To address this challenge, domain adaptation aims to bridge the gap between source and target datasets, thereby reducing the impact of dataset shift. In addition to dataset shift, EEG data often have varying numbers and positions of electrodes in different recording setups, resulting in heterogeneous data.
@@ II EEG covariance and Riemannian geometry
In this section, we introduce the Riemannian geometry concepts used for EEG covariance-based classification and transfer learning in BCIs.
@@ II-A Covariance matrix representation
As we deal with interpolated signals filled with linear combinations of the sensors’ signals, we regularised the empirical covariance matrices with the Ledoit-Wolf shrinkage to avoid rank deficiency:
$\hat{\textbf{C}}=(1-s)\textbf{C}+s\mu\mathbf{I}_{P}$ (2)
with $\mu=\operatorname{Tr}(\textbf{C})/P$ and $s$ the shrinkage coefficient [11]. The regularised covariance matrix belongs to the set $\mathcal{S}^{++}_{P}$ of Symmetric Positive Definite (SPD) matrices.
@@ III Matching EEG data dimensions
This work focuses on combining several heterogeneous datasets with varying numbers of channels due to different recording devices. Figure 1 illustrates the variability in positions of the sensors on the scalp across the six datasets considered in the experimental evaluation. In this section, we present existing strategies to deal with heterogeneous dimensionalities and propose a new source free approach with channel interpolation.
@@ III-D Physics-informed approach: Interpolation
Interpolation is a technique used in the context of EEG data processing to reconstruct the signals of malfunctioning or too noisy channels, usually referred to as ‘bad’ channels. It uses the signals from the functional channels around the bad ones. In this work, we used interpolation to map the different channels of EEG datasets: the EEG signal was reconstructed on fixed final positions based on the existing signal from all sensors of the datasets. Interpolation involves constructing a linear operator $\textbf{A}\in\mathbb{R}^{P\times P_{j}}$ that maps the $P_{j}$ existing EEG channels to the $P$ positions of a fixed template:
$\hat{\textbf{X}}=\textbf{AX}\enspace,$ (10)
where $\textbf{X}\in\mathbb{R}^{P_{j}\times T}$ are the recorded EEG signals and $\hat{\textbf{X}}\in\mathbb{R}^{P\times T}$ are the reconstructed signals. This operator can be estimated to reconstruct the EEG signal at any desired position, even if there is no corresponding sensor at that location. In the following, we present two common interpolation techniques used with EEG data: the spherical spline interpolation and the field interpolation.
III-D1 Spherical spline interpolation (SSI)
@@ IV-B Preprocessing and classification pipeline
The six datasets we used have only one channel in common: the Cz channel, located at the top of the head. The common channel selection thus resulted in keeping the signal from the Cz channel, and the associated covariance matrices came down to one value, the variance of this channel. As the union of the channels across all datasets represents 84 channels, the time series were expanded to 84 channels after ComImp, and the expanded covariance matrices after DT were of size $84\times 84$. For both interpolation methods, we set the final positions to which all epochs across all datasets were interpolated to the following 17 channels: [Fp1, Fp2, F7, F3, Fz, F4, F8, C3, Cz, C4, P3, Pz, P4, T3, T4, T5, T6]. The regularization term of the SSI was set to $1\mathrm{e}{-7}$ and the one of the FI to $1\mathrm{e}{-3}$. For all methods, we used as classifier a logistic regression from Scikit-Learn with a L2 regularization set to $C=1$.
In addition, we employed a subject-specific calibration procedure to estimate the upper bound of achievable performance. Here we use a fraction of the target subject’s data as part of the training set. This was achieved by splitting the target subject’s data into two halves, utilizing the first half for training and the second for evaluation. Consequently, this yielded a single calibration accuracy value for each subject.
@@ IV-D Results
Results for the three datasets with the most subjects are shown in Figure 3. The first row (A) illustrates the classification accuracy difference of each subject of the target dataset as a function of the number of target channels present in the training datasets. The difference was made by subtracting the FI accuracy to other methods, subject per subject. For BNCI2014_001 and PhysionetMI, ComImp and DT reach lower accuracies than FI with few target channels seen during training. When more target channels are seen in train, source free methods perform similarly to source dependent methods, with accuracies comparable to those of the calibration (purple box on row (B). For Shin2017A, FI consistently outperforms all other methods, especially ComImp, DT and common channel selection. The reason is that Shin2017A shares very few channels with the other datasets, and these channels are not located near the motor cortex. The boxplots of the second row (B) correspond to the last point of the learning curves, for which five of the datasets were used as training set. The methods are sorted according to whether they require access to target labels (supervised), source data (source dependent), or nothing (source free). For BNCI2014_001 and PhysionetMI, interpolation methods performed similarly to source dependent methods and calibration. For Shin2017A, the observations are the same as on the learning curve. Out of all six datasets, the FI led to accuracies significantly higher than the SSI for four datasets with p-values $p\leq 5\mathrm{e}{-2}$. In addition, interpolation methods were faster to compute than ComImp and DT due to the smaller size of the dimension matched data (17 channels for interpolation compared to 84 for ComImp and DT). Overall, it is important to note that FI achieved similar or better accuracies than ComImp and DT with less information.

### v2

@@ I Introduction
While machine learning is a promising tool for large-scale EEG analysis, its effectiveness can be compromised by dataset shifts. This problem occurs when a model is applied to unseen data (target) with a different distribution than the data on which it was trained (source) [5]. In EEG data, this variability arises from inter-subject differences, inter-session variability, and differences in recording equipment, e.g., amplifiers and electrode nets. To address this challenge, domain adaptation aims to bridge the gap between source and target datasets, thereby reducing the impact of dataset shift [6, 7]. In addition to dataset shift, EEG data often have varying numbers and positions of electrodes in different recording setups, resulting in heterogeneous data.
@@ II-A Covariance matrix representation
As we deal with interpolated signals filled with linear combinations of the sensors’ signals, we regularised the empirical covariance matrices with the Ledoit-Wolf shrinkage to avoid rank deficiency: $\hat{\textbf{C}}=(1-s)\textbf{C}+s\mu\mathbf{I}_{P}$ with $\mu=\operatorname{Tr}(\textbf{C})/P$ and $s$ the shrinkage coefficient [13]. The regularised covariance matrix belongs to the set $\mathcal{S}^{++}_{P}$ of Symmetric Positive Definite (SPD) matrices.
@@ III Matching EEG data dimensions
This work focuses on combining several heterogeneous datasets with varying numbers of channels due to different recording devices. Figure 1 illustrates the variability in positions of the sensors on the scalp across the six datasets considered in the experimental evaluation. We present existing strategies to deal with heterogeneous dimensionalities and propose a new source free approach with channel interpolation.
@@ III-D Physics-informed approach: Interpolation
Interpolation is a technique used in the context of EEG data processing to reconstruct the signals of malfunctioning or too noisy channels, usually referred to as ‘bad’ channels. It uses the signals from the functional channels around the bad ones. In this work, we used interpolation to map the different channels of EEG datasets: the EEG signal was reconstructed on fixed final positions based on the existing signal from all sensors of the datasets. Interpolation involves constructing a linear operator $\textbf{A}\in\mathbb{R}^{P\times P_{j}}$ that maps the $P_{j}$ existing EEG channels to the $P$ positions of a fixed template: $\hat{\textbf{X}}=\textbf{AX}$. $\textbf{X}\in\mathbb{R}^{P_{j}\times T}$ are the recorded EEG signals and $\hat{\textbf{X}}\in\mathbb{R}^{P\times T}$ are the reconstructed signals. This operator can be estimated to reconstruct the EEG signal at any desired position, even if there is no corresponding sensor at that location. Depending on the EEG montage, $P$ can either be smaller or larger than $P_{j}$. We present two interpolation techniques used with EEG data: the spherical spline interpolation (SSI) and the field interpolation (FI).
SSI
@@ IV-B Preprocessing and classification pipeline
The six datasets we used have only one channel in common: the Cz channel, located at the top of the head. The common channel selection thus resulted in keeping the signal from the Cz channel, and the associated covariance matrices came down to one value, the variance of this channel. As the union of the channels across all datasets represents $P_{\text{exp}}=84$ channels, the time series were expanded to 84 channels after ComImp, and the expanded covariance matrices after DT were of size $84\times 84$. For both interpolation methods, we set the final positions to which all epochs across all datasets were interpolated to the following $P=17$ channels: [Fp1, Fp2, F7, F3, Fz, F4, F8, C3, Cz, C4, P3, Pz, P4, T3, T4, T5, T6]. Even though we chose $P<P_{\text{exp}}$, all channels of each dataset were used to reconstruct the interpolated signals. We determined this number of final positions to ensure that the task-related data was present in the reconstructed signals. The regularization term of the SSI was set to $1\mathrm{e}{-7}$ and the one of the FI to $1\mathrm{e}{-3}$. For all methods, we used as classifier a logistic regression from Scikit-Learn with a L2 regularization set to $C=1$. In addition, we employed a subject-specific calibration procedure to estimate the upper bound of achievable performance. Here we use a fraction of the target subject’s data as part of the training set. This was achieved by splitting the target subject’s data into two halves, utilizing the first half for training and the second for evaluation. Consequently, this yielded a single calibration accuracy value for each subject.
@@ IV-D Results
Results for the three datasets with the most subjects are shown in Figure 3. The first row (A) illustrates the classification accuracy difference of each subject of the target dataset as a function of the number of target channels present in the training datasets. The difference was made by subtracting the FI accuracy to other methods, subject per subject. For BNCI2014_001 and PhysionetMI, ComImp and DT reach lower accuracies than FI with few target channels seen during training. When more target channels are seen in train, source free methods perform similarly to source dependent methods, with accuracies comparable to those of the calibration (purple box on row (B)). For Shin2017A, FI consistently outperforms all other methods, especially ComImp, DT and common channel selection. The reason is that Shin2017A shares very few channels with the other datasets, and these channels are not located near the motor cortex. The boxplots of the second row (B) correspond to the last point of the learning curves, for which five of the datasets were used as training set. The methods are sorted according to whether they require access to target labels (supervised), source data (source dependent), or nothing (source free). For BNCI2014_001 and PhysionetMI, interpolation methods performed similarly to source dependent methods and calibration. For Shin2017A, the observations are the same as on the learning curve. Out of all six datasets, the FI led to accuracies significantly higher than the SSI for four datasets with p-values $p\leq 5\mathrm{e}{-2}$. In addition, interpolation methods were faster to compute than ComImp and DT due to the smaller size of the dimension matched data (17 channels for interpolation compared to 84 for ComImp and DT). Overall, it is important to note that FI achieved similar or better accuracies than ComImp and DT with less information.